# Domain-Specific Code Language Models: Unraveling the Potential for HPC Codes and Tasks

### v1

@@ VI Tokompiler: Preprocessing to Evaluate Code Structure Understanding
As detailed in section III-section IV, it is crucial that an effective code LM for HPC tasks understands both the syntax and structure of the code, without relying on the memorization of potentially misleading human semantics [53] (e.g., variable names). We propose preprocessing code based on parsing the abstract syntax tree (AST) to remove semantic information. A code LM that performs well on this anonymized preprocessed code is clearly relying on its understanding of code structure, rather than natural language-derived semantics. We name this preprocessing method Tokompiler, inspired by input tokenization and by ensuring functionally correct and compilable code.
@@
Tokenizing code for LLMs necessitates specialized techniques to accommodate programming language syntax. [^5] These approaches include utilizing BPE and subword tokenization akin to natural language [54], employing syntax-aware tokenization to identify language-specific elements like keywords and identifiers [55], constructing tokens based on the AST to capture structural information [56], implementing language-specific lexers following grammar rules [57], preserving character integrity with character-level tokenization [58], tailoring tokenization to unique syntax rules, and leveraging dedicated code tokenization libraries. LLMs geared towards code comprehension, such as GPT-3.5-Turbo for code, likely combine several techniques, prioritizing syntax-aware tokenization to effectively process and generate code snippets in various programming languages and tasks [59]. An examination of tokenization strategies in GPT models reveals some deficits. Figure 4 presents an example of GPT3 (3(a)), Codex (3(b)), and GPT-3.5&4 (3(c)) tokenization of the Pi C code. In GPT3, one of the issues is that each individual whitespace is a new token. An advanced pre-processor is used for Codex and GPT-3.5&4. However, it is still unaware of the need to keep language primitives and numbers intact; instead, it splits numbers into subtokens.
@@
Tokompiler preprocessing (Figure 5) replaces variable names, numbers, and strings with random variable names and removes superfluous input (e.g., comments, extra whitespace). In detail:
@@
  2. 2.
    AST Generation: Parse the anonymized code using TreeSitter [^6] https://github.com/tree-sitter/tree-sitter or any suitable parser to generate an AST.
@@
    AST to Code: Transform the updated AST back into code, thus eliminating any comments that may interfere with anonymization. Although the removal of natural language (NL) from code may hamper the model’s ability to solve NL tasks, compilation tasks do not require NL understanding.
@@
    Random Number Attachment: For recurrent tokens (e.g., var_1 or num_2), use statistics to attach random integers from a predefined range (e.g., 1 to 1000) during tokenization. The attached numbers are randomly chosen without any relation to the type or order of the replaced tokens or the file/function length. This step also eliminates misleading semantics. For instance, if variable i is consistently replaced with var_1, then the model may learn that var_1 is an index variable of for loops.
VII MonoCoder Evaluation
VII-A Language Modeling Evaluation
{tikzpicture}
(a) Model Size
{tikzpicture}
(b) Perplexity C
{tikzpicture}
(c) Perplexity C++
Figure 6: Perplexity comparison across code LMs. MonoCoder was evaluated on the code from the test set of HPCorpus. The data from the other models are from [27] and [33].
Metric: Perplexity. Perplexity is a metric commonly employed in natural language processing and language modeling to assess the efficacy of a probabilistic model in predicting a given sequence of tokens. It serves as a measure of the model’s uncertainty or confusion when assigning probabilities to a set of observed data. A lower perplexity value signifies the better predictive performance of a model. In the context of our research, we measured perplexity to evaluate the predictive capabilities of the pre-trained MonoCoder language model. That is, this assessed how well the model assigned probabilities to the token sequences in the C and C++ portions of the HPCorpus test set. We then compared this perplexity to various pre-trained LLMs assessed on other C and C++ datasets in published work [27, 33]. The StarCoder values are for the model with a 2K context window, similar to MonoCoder’s 2048 context window.
Setup. For this test, we use the training procedure outlined earlier to pre-train MonoCoder on HPCorpus. We then measure perplexity on C and C++ code from the test set of HPCorpus. We compare these values to published perplexity measurements of other, larger pre-trained code LMs from other C and C++ test data.
Results. The perplexity comparison can be seen in Figure 6. We can see that despite the much smaller size of MonoCoder, it only performs slightly worse than much larger models. On C language, it is performing better than the 3x-larger GPT-Neo model.
VII-B Code Generation
In this second evaluation, we assess the ability of MonoCoder and other LLMs in generating the remainder of a code block when provided with varying amounts of prior context. While perplexity provides information about the uncertainty of the model, it does not necessarily measure the quality of the generated code. For this reason, we further evaluate code understanding with the CodeBLEU [60] score.
Metric: CodeBLEU score. In the realm of code synthesis, the selection of an appropriate evaluation metric plays a pivotal role in discerning the efficacy of machine-generated code. Commonly utilized metrics, such as BLEU [61], while effective in natural language contexts, fall short in comprehensively evaluating code quality due to their inability to account for syntactic and semantic intricacies inherent in programming languages. To address these limitations, we have chosen to employ the CodeBLEU [60] metric in our evaluation framework. CodeBLEU amalgamates the robustness of BLEU, incorporating n-gram matching, with an innovative integration of code syntax and semantics through AST and data-flow structures. This holistic approach provides a nuanced evaluation that extends beyond token matching, considering the significance of keywords, syntactic accuracy, and semantic correctness. In the case of HPC code without natural language, CodeBLEU is the appropriate metric to check similarity to generations.
Setup. For this evaluation, we first devise two sub-datasets: (1) General dataset of about 20K examples of C and C++ programs from HPCorpus with HPC-orientation (ones that used to contain parallel for pragmas. However, those pragmas were intentionally removed to keep clean C and C++ programs that have HPC orientation) and (2) OpenMP dataset of about 20K examples of C and C++ programs containing OpenMP code (e.g., parallel for). The method to test the model’s understanding is by incrementally supplying it parts of those programs (100, 300, and 600 first tokens, out of an average of 1200 tokens per code) and evaluating the outcome of the rest of the function using CodeBLEU. Demonstration of the idea is presented in Figure 7.
@@ VII-B Code Generation
2// COMPLETION:
3int i;
4int n = argc - 1;
5int m = argc - 2;
6int k = argc - 3;
7int l = argc - 4;
8int r = argc - 5;
9int c = argc - 6;
10int d = argc - 7;
11int e = argc - 8;
12int f = argc - 9;
13int g = argc - 10;
14int h = argc - 11;
15int i1 = argc - 12;
16int i2 = argc - 13;
17int i3 = argc - 14;
@@
10// COMPLETION:
11start = omp_get_wtime();
12for (i=0; i<steps; i++) {
13x = (double) i;
14pi = 3.141592653589793238
154626433832795028841971693
169937510582097494459230781
176406286208998628034825342...
@@
Figure 7: Testing generation of HPC code by the foundation models (PolyCoder example): Evaluating machine-generated text given different contexts of the initial HPC codes and measures the similarity to the true reference.
Results. 7(a) shows code completion performance of the models on General dataset. For a context of the first 100 tokens, MonoCoder achieves a CodeBLEU score of 0.62, while PolyCoder and GPT-3.5 attain scores of 0.45 and 0.44, respectively. This discrepancy can be attributed to PolyCoder’s diverse training across programming languages and GPT-3.5’s general nature, resulting in versatile but less HPC-specific answers (the limitations highlighted in section II). Expanding the context window to 600 tokens further widens the CodeBLEU score gap between MonoCoder and GPT-3.5, emphasizing lack of HPC knowledge in GPT. Overall, the model’s reduced orientation towards HPC codes accentuates the discernible gap in HPC code knowledge among the models.
{tikzpicture}
@@
{tikzpicture}
@@
Figure 8: Performance of various models on Code Completion task.
@@
Similar trends are evident on OpenMP dataset in 7(b), depicting the CodeBLEU scores of models when completing functions containing OpenMP pragmas. Generally, the results are slightly lower than the those on General dataset, underscoring the models’ limited understanding of OpenMP.
@@
It is noteworthy that, despite the Tokompiler pre-processing having a minor impact on the performance of MonoCoder and GPT-3.5, this code representation significantly impairs the performance of PolyCoder. This observation suggests that PolyCoder heavily relies on local semantics when completing programs.
VIII Downstream HPC Task: OpenMP Pragma Generation
In this section, we evaluate the ability of MonoCoder and other existing LLMs in assisting on HPC specific downstream task. Specifically, we evaluate ability of these models in automatically suggesting appropriate OpenMP pragma for the input C/C++ program.
Task objective. While a lot of progress has been made in integrating machine learning into the identification of parallelism opportunities in C/C++ programs [15, 26, 62], there has been a notable gap in connecting parallelism detection with code generation. In an effort to bridge this gap, our work leverages MonoCoder, refining its capabilities through fine-tuning on the OpenMP-specific subcorpus of HPCorpus. Specifically, we created OMPify+ by fine-tuning MonoCoder to generate complete OpenMP pragmas, when presented with a for loop written in C/C++ languages. OpenMP pragma annotations on for loops allow such loops to improve performance by using multi-threading to process computations inside the loop concurrently.
Dataset. To create a dataset for automatic OpenMP parallelization task, we curated a subcorpus of HPCorpus named HPCorpusOMP . This subcorpus specifically includes all the for loops that have either #omp parallel for pragma that distributes the workload across multiple CPU cores, or offload tasks to a team of threads for GPU execution using #omp target teams distribute. The dataset also includes OpenMP pragmas containing private and reduction clauses. The details of the dataset in terms of these clauses is present in Table II.
private reduction target Total
C 3,526 713 145 8,764
C++ 2,122 1,424 345 18,151
TABLE II: Clause breakdown in HPCorpusOMP .
It is important to note that we excluded context of the loops to maintain focus on the loop-specific information. However, to avoid the loss of crucial details, we designed HPCorpusOMP such that the private clause encompasses instances of both firstprivate and lastprivate, as both of these clauses serve the purpose of creating a private instance of variables for each thread.
Evaluation setup. We finetuned MonoCoder on the HPCorpusOMP subcorpus. The finetuning task was setup as a generation task in which the complete for loop is input to the model, and the expected generated output was the pragma and its associated variables. We then evaluated MonoCoder on a test set of HPCorpusOMP , both with the original for loops and their anonymized versions using Tokompiler. For comparison, we assessed GPT-3.5 turbo in a zero-shot manner. Specifically, we supplied a prompt instructing it to “Generate the optimal OpenMP pragma for the provided code” when presented with the for loop.
Fine-tuning details. We fine-tuned the pretrained MonoCoder model to the OpenMP pragma generation task using the Huggingface transformers library on 2 NVIDIA V100 32GB GPUs at fp32 precision. Fine-tuning used the AdamW optimizer with a linear warmup over the first 100 steps, followed by a linear decay over the remaining steps. Fine-tuning samples also had a maximum length of 2048 tokens, trained in 16 sample minibatches (8 per GPU). The fine-tuning was run for 4 epochs at the learning rate of 0.000016.
(a) Clause evaluation.
(b) Inner-clause evaluation.
Figure 9: Classification tasks for pragma evaluation.
{tikzpicture}
(a) Performance on classifying private clause (Task \@slowromancapi@).
{tikzpicture}
(b) Performance on classifying reduction clause (Task \@slowromancapii@).
Figure 10: Performance of MonoCoder and GPT on OpenMP pragma classification tasks.
Figure 11: Generation of private clause (Task \@slowromancapiii@).
{tikzpicture} (a) Accuracy of predicting full private clause. {tikzpicture} (b) Cumulative distribution of samples.
{tikzpicture} (a) Accuracy of predicting full reduction clause. {tikzpicture} (b) Cumulative distribution of samples.
Figure 11: Generation of private clause (Task \@slowromancapiii@).
Figure 12: Generation of reduction clause (Task \@slowromancapiv@).
Evaluation method. While we used the CodeBLEU score to evaluate the fluency of generated code containing OpenMP pragmas, it falls short in capturing correctness aspects. To comprehensively evaluate the performance of MonoCoder and GPT, we broke down the performance measurement of the OpenMP pragma generation task into five subtasks:
  1. 1.
    Clause-level generation: This set of tasks evaluates whether the model correctly generates OpenMP private (Task \@slowromancapi@) and reduction (Task \@slowromancapii@) clauses. For example, in the sample depicted in 8(a), the evaluation involves marking true-positive (TP) for the private clause as it is present in both the label and prediction. Conversely, for the reduction clause, we mark it as false-negative (FN) as the model fails to generate the required clause present in the label. We also report precision and recall to account for true negatives (TN), since many samples do not contain private or reduction clauses. For the reduction operator, we will focus solely on accuracy, since the result is deemed correct if the operator is the same and incorrect otherwise. The results below depict the outcomes of these evaluations.
  2. 2.
    Generating the correct variables of OpenMP clauses: This set of tasks evaluates the correctness of the generated variables for each private (Task \@slowromancapiii@) and reduction clause (Task \@slowromancapiv@). For the reduction clause, we also measure the accuracy of generating the reduction operator (Task \@slowromancapv@). Our evaluation recognizes that clause variables, like clauses themselves, could be unordered. 8(b) illustrates the evaluation method with an example. For example, variables A and B would be considered true-positive (TP), while variable X would be a false-positive (FP), as it is present in the generated code but not in the ground-truth label.
Results on clause-level generation. In Task \@slowromancapi@, MonoCoder provided accuracy of 94.4%, surpassing GPT 3.5, which achieved 77% (9(a)). The substantial gap in recall between MonoCoder and GPT can attributed to GPT frequently mispredicting the need for the private clause (FN), thereby impacting its overall performance. Similarly, in Task \@slowromancapii@, MonoCoder exhibited an accuracy of 98.3%, outperforming GPT with an accuracy of 92.3% (9(b)).
An intriguing observation is the minimal impact of the anonymized Tokompiler version on MonoCoder’s results, indicating robustness. In contrast, GPT’s performance significantly degraded with the anonymized version, suggesting its reliance on local semantics rather than the underlying code structure, unlike MonoCoder. These findings emphasize MonoCoder’s ability to leverage code structure for accurate predictions in contrast to GPT, which heavily depends on local semantic information.
Results on generating clause variables. OpenMP private and reduction clauses can contain different number of variables. To understand the impact of different number of variables, we report performance of Task \@slowromancapiii@ and \@slowromancapiv@ in terms of the number of variables in those clauses.
In Task \@slowromancapiii@ (Figure 12), the analysis reveals that the majority of private clauses involve up to five variables. For private clauses with one variable, MonoCoder generates 69% of the variables, while GPT succeeds only with 5.3%. When examining all the clauses, regardless of their variable count, MonoCoder accurately generates 48.5% of these variables, while GPT scores only 14.3%. We can make two key observations from these results. Firstly, for codes that do not include many potential variables that should be private, it is easier to identify the variables that should be private to each thread. Furthermore, the substantial performance gap between GPT and MonoCoder in this examination underscores a significant issue: GPT often fails to generate variables within the clause, even when they are present in the labeled clause.
Similar trends are observed in Task \@slowromancapiv@ (Figure 12). For reduction clauses with a single variable, MonoCoder achieves 58.1% accuracy, while GPT achieves 55.2%. When examining all the reduction clauses, MonoCoder attains an accuracy of 51.7%, while GPT achieves 51.2%. Notably, the anonymized Tokompiler version has a minimal impact on MonoCoder’s variable generation. In stark contrast, the anonymized version significantly degrades GPT’s performance, underscoring its deficiency in comprehending code structure.
In Task \@slowromancapv@, MonoCoder generates the reduction operator correctly 81% of the time, while GPT succeeds 79% of the time. The similar performance between GPT and MonoCoder, along with the minimal impact of the anonymized Tokompiler version on both the models, can be explained by the inherent simplicity of the task. The detection of the reduction operator becomes relatively straightforward when there is a variable aggregating the computation along the loop iterations.
IX Related Work
The code LM literature is somewhat divided into NLP-oriented approaches and software engineering approaches. The NLP-oriented approaches typically evaluate code LMs with perplexity and datasets like HumanEval [50] or Mostly Basic Programming Problems [63], which assess both natural language comprehension and general reasoning ability along with code comprehension. Evaluations more specific to code include tasks such as code search or code repair [64]. Only very recently have we seen an evaluation of code LMs for HPC-specific tasks, such as OpenMP pragma prediction [16] and parallel code generation [29], and race detection [28]. Our work complements these recent papers by demonstrating how a smaller, domain-specific LM can accomplish HPC tasks. We also demonstrate a novel way to assess how heavily a model relies on local semantics, rather than a structural understanding of the code, to perform various tasks such as code completion.
While we have mentioned work on domain-specific LMs [30, 31, 32] in section II, there are also related investigations in NLP on improving the reasoning and few-shot learning capabilities of smaller LMs [65, 66, 67]. These investigations designed procedures tailored to the task of interest to enable smaller models to perform as well, if not better than, larger ones. Our work contributes to a general trend in improving LMs for use in low-resource settings.
X Conclusion
Large language models, such as GPT-3.5, GPT-4, etc., have revolutionized the field of natural language processing, with their chat-based versions, such as ChatGPT, catching significant attention worldwide. This has naturally led to their applications in programming and, specifically, high-performance computing. Our experiments with existing LLMs, however, revealed that these LLMs cannot solve HPC-related programming problems accurately, or they are multi-lingual and are trained on languages that are not HPC-related. More importantly, these models are huge in size and demand expensive compute resources to train.

### v3

@@ IV Local Semantics Elimination (LSE)
An effective code LM for HPC tasks must understand both the syntax and structure of the code, without memorizing potentially misleading human semantics [36] (e.g., variable names). We propose a preprocessing method based on an abstract syntax tree (AST) to remove misleading semantic information. A code LM that performs well on this anonymized preprocessed code is clearly relying on its understanding of code structure, rather than natural language-derived semantics. We name this preprocessing method as LSE (for local semantics elimination). LSE is inspired by input tokenization and more importantly ensures functionally correct and compilable code.
@@
Tokenizing code for LLMs necessitates specialized techniques to accommodate programming language syntax. [^1] These approaches include utilizing BPE and subword tokenization akin to natural language [37], employing syntax-aware tokenization to identify language-specific elements like keywords and identifiers [38], constructing tokens based on the AST to capture structural information [39], implementing language-specific lexers following grammar rules [40], preserving character integrity with character-level tokenization [41], tailoring tokenization to unique syntax rules, and leveraging dedicated code tokenization libraries. LLMs geared towards code comprehension, such as GPT-3.5-Turbo for code, likely combine several techniques, prioritizing syntax-aware tokenization to effectively process and generate code snippets in various programming languages and tasks.
@@
In short, LSE preprocessing (Figure 1) replaces variable names, numbers, and strings with random variable names and removes superfluous input (e.g., comments, extra whitespace). The detailed steps are as follows:
@@
    AST Generation: Parse the code using TreeSitter [^2] https://github.com/tree-sitter/tree-sitter or any suitable parser to generate an AST.
  2. 2.
@@
    AST to Code: Transform the updated AST back into code, while eliminating any comments that may interfere with anonymization.
@@
    Random Number Attachment: For recurrent tokens (e.g., var_1 or num_2), attach random integers from a predefined range (e.g., 1 to 1000) during tokenization. The attached numbers are randomly chosen without any relation to the type or order of the replaced tokens or the file/function length. This step also eliminates misleading semantics. For instance, if variable i is consistently replaced with var_1, then the model may learn that var_1 is an index variable of for loops.
// INPUT to LLM:
@@
// LLM COMPLETION:
2int i;
3int n = argc - 1;
4int m = argc - 2;
5int k = argc - 3;
6int l = argc - 4;
7int r = argc - 5;
8int c = argc - 6;
9int d = argc - 7;
10int e = argc - 8;
11int f = argc - 9;
12int g = argc - 10;
13int h = argc - 11;
14int i1 = argc - 12;
15int i2 = argc - 13;
@@
// LLM COMPLETION:
10start = omp_get_wtime();
11for (i=0; i<steps; i++) {
12x = (double) i;
13pi = 3.141592653589793238
144626433832795028841971693
159937510582097494459230781
166406286208998628034825342...
@@
Figure 3: Evaluating code completion performance of HPC code by the foundation models (PolyCoder example): Evaluating machine-generated code, given different contexts of the initial HPC codes and measuring the similarity to the true reference.
context-100context-300context-600000.50.50.50.511111.51.51.51.50.620.620.620.620.750.750.750.750.850.850.850.850.640.640.640.640.720.720.720.720.850.850.850.850.450.450.450.450.640.640.640.640.780.780.780.780.390.390.390.390.60.60.60.60.740.740.740.740.440.440.440.440.580.580.580.580.590.590.590.590.530.530.530.530.580.580.580.580.60.60.60.6CodeBLEUMonoCoderMonoCoder + LSEPolyCoderPolyCoder + LSEGPT-3.5GPT-3.5 + LSE
@@
context-100context-300context-600000.50.50.50.511111.51.51.51.50.640.640.640.640.770.770.770.770.840.840.840.840.640.640.640.640.770.770.770.770.860.860.860.860.610.610.610.610.740.740.740.740.840.840.840.840.370.370.370.370.590.590.590.590.740.740.740.740.430.430.430.430.570.570.570.570.560.560.560.560.520.520.520.520.570.570.570.570.60.60.60.6CodeBLEUMonoCoderMonoCoder + LSEPolyCoderPolyCoder + LSEGPT-3.5GPT-3.5 + LSE
@@
Figure 4: Code Completion Performance on General and OpenMP Datasets –- CodeBLEU scores (higher is better) for MonoCoder, PolyCoder, and GPT-3.5 models, both with and without Local Semantic Elimination (LSE), across varying context lengths (100, 300, and 600 tokens). MonoCoder and MonoCoder + LSE consistently outperform other models, with the addition of LSE generally enhancing performance across all models.
V MonoCoder Evaluation
V-A Language Modeling Evaluation
Metric: Perplexity. Perplexity is a metric commonly employed in natural language processing and language modeling to assess the efficacy of a probabilistic model in predicting a given sequence of tokens. It serves as a measure of the model’s uncertainty or confusion when assigning probabilities to a set of observed data. A lower perplexity value signifies the better predictive performance of a model. In the context of our research, we measured perplexity to evaluate the predictive capabilities of the pre-trained MonoCoder language model. That is, this assessed how well the model assigned probabilities to the token sequences in the C and C++ portions of the HPCorpus test set.
Setup. For this test, we use the training procedure outlined earlier to pre-train MonoCoder on HPCorpus. We then measure perplexity on C and C++ code from the test set of HPCorpus. We then compared this perplexity to the perplexity scores of various pre-trained LLMs based on the published works [24, 33]. The StarCoder values are with a 2K context window, similar to MonoCoder’s 2048 context window.
Results. The perplexity comparison can be seen in Figure 2. We can see that despite the much smaller size of MonoCoder, it suffers from minor performance degradation than much larger models. For C language, it is performing better than the 3x-larger GPT-Neo model.
V-B Code Completion Evaluation
In this second evaluation, we assess the ability of MonoCoder and other LLMs in completing a block of code when provided with varying amounts of prior context. While perplexity provides information about the uncertainty of the model, it does not necessarily measure the quality of the generated code. For this reason, we further evaluate code understanding with the CodeBLEU [42] score.
Metric: CodeBLEU score. CodeBLEU [42] is the metric of choice in the code completion tasks. CodeBLEU amalgamates the robustness of BLEU score [43](from NLP), incorporating n-gram matching, with an innovative integration of code syntax and semantics through AST and data-flow structures. This holistic approach provides a nuanced evaluation that extends beyond token matching, considering the significance of keywords, syntactic accuracy, and semantic correctness. In the case of HPC code without natural language, CodeBLEU is the appropriate metric to check similarity to generations.
Setup. For this evaluation, we first devise two sub-datasets: (1) General dataset of about 20k examples of C and C++ programs from HPCorpus with HPC-orientation [^3] Ones that contained OpenMP’s parallel for pragmas. However, those pragmas were intentionally removed to keep clean C and C++ programs that have HPC orientation and (2) OpenMP dataset of about 20k examples of C and C++ programs containing OpenMP code (e.g., parallel for). The method to test the model’s understanding is by incrementally supplying it more parts of those programs (first 100, 300, and 600 tokens, out of an average of 1200 tokens per code) and comparing the code generated by the model with the expected ground-truth code using CodeBLEU. Demonstration of the idea is presented in Figure 3.
Results. 4(a) shows code completion performance of the models on General dataset. For a context of the first 100 tokens, MonoCoder achieves a CodeBLEU score of 0.62, while PolyCoder and GPT-3.5 attain scores of 0.45 and 0.44, respectively. This discrepancy can be attributed to PolyCoder’s diverse training across programming languages and GPT-3.5’s general nature, resulting in versatile but less HPC-specific answers. Expanding the context window to 600 tokens further widens the CodeBLEU score gap between MonoCoder and GPT-3.5, emphasizing lack of HPC knowledge in GPT. Overall, the model’s reduced orientation towards HPC codes accentuates the discernible gap in HPC code knowledge among the models.
@@ V-B Code Completion Evaluation
Similar trends are evident on OpenMP dataset in 4(b), depicting the CodeBLEU scores of models when completing functions containing OpenMP pragmas. Generally, the results are slightly lower than those on General dataset, underscoring the models’ limited understanding of OpenMP.
@@
It is noteworthy that, despite the LSE pre-processing having a minor impact on the performance of MonoCoder and GPT-3.5, this code representation significantly impairs the performance of PolyCoder. This observation suggests that PolyCoder heavily relies on local semantics for code completion.
VI Related Work
We present the related work along two directions: 1) code LMs that are not specifically designed for HPC tasks but can solve HPC tasks, 2) LLMs designed specifically for HPC tasks.
The code LM literature is somewhat divided into NLP-oriented approaches and software engineering approaches. Popular LLMs such as GPT-3.5, GPT-4, LLaMa, or even code specific LLMs such as CodeT5 [44], etc., are trained on code datasets that also contain natural language comments. As a result, most of these models, if not all, can solve programming tasks with or without natural language prompts. These NLP-oriented approaches typically evaluate code LMs with perplexity and datasets like HumanEval [32] or Mostly Basic Programming Problems [45], which assess both natural language comprehension and general reasoning ability of a model along with code comprehension. These models also solve coding-specific tasks such as code search, code completion, etc. In addition to natural language and programming-specific tasks, these models can also solve HPC tasks to certain extent. Because the training datasets of these models most likely contain HPC code also (in APIs such as OpenMP, CUDA, etc.), these models possess some understanding of HPC languages also. Our experiments with GPT-3.5 revealed that it had a reasonable understanding of code parallelization problem (e.g., analyzing if a loop can be parallelized). However, our experiments also revealed that these LLMs have limited understanding of HPC tasks — the same observation also reported by Nichols et al. in their evaluation of LLMs for HPC tasks [23]. Specifically, their experimental results reveal several limitations of existing LLMs for HPC tasks.
Given the limitations of popular LLMs on HPC tasks and also their expensive training costs, several works have recently explored HPC specific LMs. Specifically, several groups have developed LMs for popular HPC problems such as OpenMP pragma prediction and generation [8, 46, 10, 12, 11, 26, 13, 9], MPI code generation [14, 47], and race detection [25]. We found that these works mostly finetune a pre-trained LM, which are larger in size than required for their HPC tasks. We believe that our work complements these papers by demonstrating a smaller, domain-specific LM that can solve HPC tasks.
VII Conclusion & Future Work